Two speakers, 1 and 2, are set up with their cente
Answer #1

Sue experuence construtive interefrence

from the given data,

wavelength of the sound wave, lamda = v/f

= 343/225

= 1.524 m


path difference at the location of call, r2 - r1 = sqrt(2)*6 - 6

= 2.485 m


path difference at the location of pam, r2 - r1 = sqrt(6^2 + 5^2) - sqrt(6^2+1^2)

= 1.727 m

path difference at the location of Bob, r2 - r1 = sqrt(6^2 + 4^2) - sqrt(6^2+2^2)

= 0.886 m

path difference at the location of Sue, r2 - r1 = sqrt(6^2 + 3^2) - sqrt(6^2+3^2)

= 0 m

path difference at the location of Tom, r2 - r1 = sqrt(6^2 + 4^2) - sqrt(6^2+2^2)

= 0.886 m


for constructive interefrence path difference should be = n*lamda (n = 0,1,2,3,4,5...)

so, Sue experuence construtive interefrence